OpenStudy (anonymous):

Solve for B R=x(A+B) I am super confused, please help! (:

OpenStudy (anonymous):

@abigailsimmons

OpenStudy (bibby):

You want to get the term that contains B isolated on one side of the equation. Here we start by determining which term contains B

OpenStudy (anonymous):

What do you mean, @bibby ? I have had these types of problems before but I really don't understand this one

OpenStudy (bibby):

if we were solving for B A+B=C we'd subtract A from both sides to get B alone here we have \(R=x(A+B)\) now we have to divide both sides by x to get (A+B) alone (the term that contains B) \(\large \frac{R}{X}=A+B\)

OpenStudy (anonymous):

Ohhhh...that starts to make snese... I think i got it

OpenStudy (bibby):

x(A+B) = x* (A+B) because the A+B is in parentheses it's easier to think of it as a single entity then we can do the opposite of division to both sides. so what do you think the answer is?

OpenStudy (anonymous):

The answer would be B=(R-Ax)/x

OpenStudy (bibby):

Weird answer. what I did was \(R=x(A+B)\) divide both sides by x \(\huge \frac{R}{x}=\frac{}{\cancel{x}}\cancel{x}(A+B)\) \(\huge \frac{R}{x}-A=(B)\) subtract both sides by a

OpenStudy (bibby):

I don't see the point in getting A in terms of x, although as far as I can tell your answer is right